# San Jose/Seattle, US https://bytedance.github.io/vidi-website/

### v2

@@ 6.2 Evaluation Metric
For each sample, we compute precision $P$, recall $R$ and IoU as the core evaluation metrics. According to the definition in Equation (5), a perfect prediction would be exactly the same as the ground-truth timestamp ranges, i.e., $P=R=\text{IoU}=1$. To capture performance across varying levels of overlap, we sweep thresholds in the range $[0,1]$ to generate accuracy-threshold curves. We then compute the area under curve (AUC) for each metric, denoted as $\bar{P}$, $\bar{R}$, $\bar{\text{IoU}}$, which serve as the final evaluation scores. While AUC provides a comprehensive summary, user experience in real-world applications often depends on performance as specific thresholds, e.g., IoU@0.5. Therefore, we also report accuracy-threshold curves (see Figure 10) to facilitate a more detailed analysis, while treating AUC as the primary metric for evaluation.
@@ 7.1 Implementation Details
We implement the released Vidi model using three core components: the visual encoder google/siglip
@@
[34], the audio encoder openai/whisper-large-v3 [22], and the language model mistralai/Mistral-7B-Instruct-v0.3 [11]. Unlike existing video LLMs that use a fixed number of uniformly sampled frames, Vidi adopts modality-specific fixed sampling rates: $1.0$ fps for video frames and $16,000$ Hz for audio signals. It ensures sufficient coverage of fine-grained visual and auditory details across varying video durations. After sampling, each image frame is independently encoded by SigLip to convert it into a sequence of visual tokens. In parallel, the audio waveform is segmented into batches, each independently encoded by Whisper, and then concatenated back into a sequence of audio tokens. The complete set of training configurations and hyper-parameters for various stages is provided in Table 2.
@@
Ultra-Short ($<1m$) $\bar{P}$ 64.5 72.3 58.8 53.6
$\bar{R}$ 79.6 65.4 69.2 42.1
$\bar{\text{IoU}}$ 54.6 49.2 42.6 32.5
Short ($1-10m$) $\bar{P}$ 57.4 51.7 43.3 30.0
$\bar{R}$ 55.8 50.6 41.7 26.5
$\bar{\text{IoU}}$ 40.5 31.6 22.7 15.6
Medium ($10-30m$) $\bar{P}$ 47.4 35.1 31.1 18.2
$\bar{R}$ 46.6 34.4 22.6 21.1
$\bar{\text{IoU}}$ 32.1 18.2 9.9 9.1
Long ($30-60m$) $\bar{P}$ 38.9 26.3 29.9 18.2
$\bar{R}$ 44.9 12.9 11.9 17.7
$\bar{\text{IoU}}$ 27.6 7.1 4.9 9.2
Ultra-Long ($>60m$) $\bar{P}$ 36.7 17.1 25.3 20.4
$\bar{R}$ 46.7 8.3 5.8 19.6
$\bar{\text{IoU}}$ 27.5 2.9 2.4 9.5
Keyword $\bar{P}$ 54.4 50.3 47.9 37.4
$\bar{R}$ 57.4 33.0 33.4 26.4
$\bar{\text{IoU}}$ 39.4 21.9 17.4 18.3
Phrase $\bar{P}$ 45.3 41.4 41.7 25.1
$\bar{R}$ 48.1 36.6 31.1 23.7
$\bar{\text{IoU}}$ 32.4 22.3 16.0 13.2
Sentence $\bar{P}$ 47.5 31.1 30.2 17.5
$\bar{R}$ 52.3 34.2 23.3 22.2
$\bar{\text{IoU}}$ 34.7 19.7 12.6 10.0
Audio $\bar{P}$ 34.2 32.5 30.9 7.3
$\bar{R}$ 43.8 35.8 25.5 16.5
$\bar{\text{IoU}}$ 24.2 20.3 10.2 3.7
Vision $\bar{P}$ 53.6 48.6 48.2 41.3
$\bar{R}$ 61.6 38.9 37.9 30.5
$\bar{\text{IoU}}$ 43.9 25.4 21.8 22.0
Vision+Audio $\bar{P}$ 51.2 37.2 33.9 21.2
$\bar{R}$ 49.0 30.9 23.4 22.0
$\bar{\text{IoU}}$ 33.4 18.4 12.1 11.1
Overall $\bar{P}$ 49.0 40.3 39.3 25.9
$\bar{R}$ 52.5 34.6 28.9 24.0
$\bar{\text{IoU}}$ 35.4 21.2 15.2 13.6
Table 3: Performance of different models on the VUE-TR benchmark across various evaluation attributes. $\bar{P}$ and $\bar{R}$ represent the AUC (Area Under Curve) values for precision and recall, respectively; while $\bar{\text{IoU}}$ denotes the AUC of intersection-over-union between prediction and ground-truth timestamp ranges, as defined in Section 6.2. GPT models are accessed via the Azure API, and Gemini is accessed via internal Google API. $\dagger$ Gemini models are evaluated by directly uploading videos. To comply with the $100$ MB upload limit, long videos are resized to a resolution of $256$ pixels. Compared to Gemini-2.0-Flash, Gemini-2.5-Pro incurs higher latency and token cost for reasoning and exhibits a higher content filtering rate, often resulting in empty outputs. $\ddagger$ GPT-4o is constrained by the Azure API’s $120$-frame limit. For videos longer than $120$ seconds, we uniformly sample $120$ frames.
@@ 7.2 Temporal Retrieval Results
Since GPT-4o API does not support audio, we extract frames at $1$ fps and feed them as input images. To adapt GPT-4o for temporal retrieval, we design a simple custom prompt with in-context examples, ensuring that the output only contains frame index ranges. An example instruction is:
@@
Unlike GPT-4o, Gemini models [1] naturally support text, vision and audio with extremely long context length, making them ideal for long video understanding. Gemini can also take raw video files as input. We evaluate two versions:
  • Gemini-2.5-Pro (exp/preview): Offers strong reasoning and often outputs explicit Chain-of-Thought (CoT) explanations. However, when not constrained by output formatting instructions, the responses can be highly inconsistent and challenging to parse. Conversely, adding strict format requirements improves consistency but leads to degraded performance. Therefore, we choose the simple prompt without format requirement and parse the major possible formats for evaluation.
  • Gemini-2.0-Flash (stable): Provides better output structure and lower rejection rates, making it more reliable for batch evaluations. Although slightly less capable in reasoning than 2.5-Pro, it balances performance and robustness well.
In Table 3, we report the performance across multiple evaluation dimensions: video duration, query format, and query modality. Remarkably, Vidi outperforms all baseline models across all categories for the primary metric $\bar{\text{IoU}}$. On ultra-short videos, Gemini-2.0-Flash achieves the best precision, but its recall is significantly lower, leading to a lower $\bar{\text{IoU}}$. Performance across different query formats is relatively consistent, though models generally struggle more with longer and more descriptive queries. As expected, GPT-4o performs poorly on audio-only queries due to lack of audio input support. Other models perform comparably across modalities, but accuracy on audio-based queries is consistently lower than that on vision-based queries, highlighting the challenge of audio-conditioned understanding.
As shown in Figure 10, we can still observe that Vidi consistently outperforms all competitors by a significant margin across the entire range of thresholds. This performance gap even widens at higher thresholds (e.g., $\geq 0.5$), where other models degrade much more rapidly. It shows Vidi’s strength in fine-grained temporal precision, a critical capability for real-world video editing and retrieval tasks.
@@ 8 Related Work
In contrast, temporal retrieval requires fine-grained understanding across the full temporal span of a video instead of a few scattered glimpses. As such, temporal retrieval serves as a more rigorous benchmark for evaluating a model’s true capability in video understanding. Charades-STA [6] and DiDeMo [9] primarily focus on action grounding within short clips ($\sim 30s$). QVHighlights [14] and ActivityNet Captions [12] extend to several minutes videos, covering event descriptions and subjectively defined highlights. Recently, LongVALE [7] introduces multimodal audio-visual queries, but remains constrained to relatively short durations.
@@
Video understanding has experienced significant progress with the emergence of LMMs that integrate visual and textual information. Models such as InternLM-XComposer [35, 4, 36], Qwen-VL [2, 28, 3], and LLaVA [19, 32] series have introduced innovative architectures and training strategies. InternLM-XComposer adopts a modular architecture that integrates interleaved text-image composition and comprehension with rich multilingual knowledge. Qwen-VL and its successors introduce flexible multi-image and box-conditioned input capabilities, supporting multilingual and multi-turn interaction for diverse vision-language tasks. The LLaVA series bridge visual perception with LLMs through a unified representation space, and its recent extensions like LLaVA-NeXT further support video and 3D input via interleaved multimodal pretraining [17]. While these models demonstrate impressive general-purpose reasoning, they typically operate on a limited number of frames (e.g., $\leq 64$), which constrains their effectiveness in long-form temporal retrieval.
@@
Early methods [14, 6, 9, 12] usually employ proposal-based strategies, where models first generate candidate segments and then rank them based on their semantic relevance to the query. With the emergence of video LMMs, recent research has shifted toward enhancing video-text alignment through unified frameworks. For instance, TimeChat [24] connects time-aware frame encoder and a sliding video Q-Former [18] to deal with temporal retrieval based on an instruction-tuning dataset. TimeSearch [21] proposes a temporal spotlight grounding strategy to find key events and a temporal reflection mechanism to verify time range predictions and guide the search direction. Ye et al. [33] revisit temporal search in long-form videos by proposing a lightweight framework that reformulates temporal retrieval as a spatial search problem. Although efficient, the above methods rely on sampling only $8\sim 96$ frames, which may be insufficient to capture fine-grained temporal cues.
Despite these advancements, most existing methods still focus on short videos or rely on sparse frame sampling, and primarily limited to vision-only inputs. These constraints hinder the evaluation of models’ full multimodal reasoning abilities, particularly in understanding and retrieving moments from long, complex videos involving both visual and auditory modalities. In this work, we propose Vidi, a new approach that jointly incorporates text, vision and audio input from hour-long videos to retrieve the most relevant temporal segments based on free-form natural language queries. Our setting pushes beyond existing benchmarks by requiring fine-grained temporal understanding across extended multimodal content.
@@ 9 Conclusion
We introduce Vidi, a large multi-modal model for real-world video understanding and editing (VUE). The first release focuses on temporal retrieval (TR), which is a foundational step in video trimming and editing, by localizing relevant segments from long videos given natural language queries. To support robust evaluation under realistic conditions, we propose VUE-TR, a new benchmark for practical video editing use cases. VUE-TR introduces key improvements over prior datasets in terms of video duration, audio support, query format, annotation quality, and metric design for multi-span timestamp evaluation. Vidi leverages a unified vision-audio-language architecture with modality-aware sampling and precise token alignment. Besides, the Decomposed Attention mechanism allows the model to capture fine-grained temporal cues across diverse modalities and scales seamlessly to long videos. As demonstrated on the VUE-TR benchmark, Vidi consistently outperforms leading proprietary models such as GPT-4o and Gemini. For future work, we plan to extend Vidi to support a broader spectrum of video understanding tasks, including temporal VQA, and high-level video understanding. Additionally, we aim to explore interactive editing capabilities to further bridge the gap between large models and real-world VUE applications.
@@ 10 Contributors
Chia-Wen Kuo, Dawei Du, Fan Chen, Guang Chen, Sijie Zhu.
Core Contributors - Infrastructure (alphabetical order) s
@@
Jiamin Yuan, Lingxi Zhang, Lu Guo, Lusha Li, Qingyu Chen, Rachel Deng, Stuart Siew, Wei Lu, Wen Zhong, Xin Gu, Xing Mei, Xueqiong Qu.

### v3

@@ 6.2 Evaluation Metric
For each sample, we compute precision $P$, recall $R$ and IoU as the core evaluation metrics. According to the definition in Equation (5), a perfect prediction would be exactly the same as the ground-truth timestamp ranges, i.e., $P=R=\text{IoU}=1$. To capture performance across varying levels of overlap, we sweep thresholds in the range $[0,1]$ to generate accuracy-threshold curves. We then compute the area under curve (AUC) for each metric, denoted as $\bar{P}$, $\bar{R}$, $\bar{\text{IoU}}$, which serve as the final evaluation scores. Although AUC provides a comprehensive summary, the user experience in real-world applications often depends on performance as specific thresholds, e.g., IoU@0.5. Therefore, we also report accuracy-threshold curves (see Figure 10) to facilitate a more detailed analysis, while treating AUC as the primary metric for evaluation.
@@ 7.1 Implementation Details
We implement the released Vidi models using three core components: the visual encoder (google/siglip
@@
[35] for Vidi, and google/siglip2-so400m-patch14-384 [27] for Vidi-1.5, the audio encoder openai/whisper-large-v3 [22], and the language model ( mistralai/Mistral-7B-Instruct-v0.3 [11] for Vidi, and google/gemma-2-9b [25] for Vidi-1.5). Unlike existing video LLMs that use a fixed number of uniformly sampled frames, Vidi adopts modality-specific fixed sampling rates: $1.0$ fps for video frames and $16,000$ Hz for audio signals. It ensures sufficient coverage of fine-grained visual and auditory details across varying video durations. After sampling, each image frame is independently encoded by SigLip to convert it into a sequence of visual tokens. In parallel, the audio waveform is segmented into batches, each independently encoded by Whisper, and then concatenated back into a sequence of audio tokens. The complete set of training configurations and hyper-parameters for various stages is provided in Table 2.
@@
Ultra-Short ($<1m$) $\bar{P}$ 70.7 64.5 72.3 58.8 53.6
$\bar{R}$ 76.2 79.6 65.4 69.2 42.1
$\bar{\text{IoU}}$ 58.4 54.6 49.2 42.6 32.5
Short ($1-10m$) $\bar{P}$ 60.5 57.4 51.7 43.3 30.0
$\bar{R}$ 63.4 55.8 50.6 41.7 26.5
$\bar{\text{IoU}}$ 44.1 40.5 31.6 22.7 15.6
Medium ($10-30m$) $\bar{P}$ 53.8 47.4 35.1 31.1 18.2
$\bar{R}$ 54.0 46.6 34.4 22.6 21.1
$\bar{\text{IoU}}$ 38.1 32.1 18.2 9.9 9.1
Long ($30-60m$) $\bar{P}$ 47.8 38.9 26.3 29.9 18.2
$\bar{R}$ 40.6 44.9 12.9 11.9 17.7
$\bar{\text{IoU}}$ 30.5 27.6 7.1 4.9 9.2
Ultra-Long ($>60m$) $\bar{P}$ 46.1 36.7 17.1 25.3 20.4
$\bar{R}$ 43.2 46.7 8.3 5.8 19.6
$\bar{\text{IoU}}$ 32.3 27.5 2.9 2.4 9.5
Keyword $\bar{P}$ 60.5 54.4 50.3 47.9 37.4
$\bar{R}$ 61.5 57.4 33.0 33.4 26.4
$\bar{\text{IoU}}$ 43.8 39.4 21.9 17.4 18.3
Phrase $\bar{P}$ 53.1 45.3 41.4 41.7 25.1
$\bar{R}$ 50.2 48.1 36.6 31.1 23.7
$\bar{\text{IoU}}$ 36.8 32.4 22.3 16.0 13.2
Sentence $\bar{P}$ 52.9 47.5 31.1 30.2 17.5
$\bar{R}$ 53.1 52.3 34.2 23.3 22.2
$\bar{\text{IoU}}$ 38.5 34.7 19.7 12.6 10.0
Audio $\bar{P}$ 38.1 34.2 32.5 30.9 7.3
$\bar{R}$ 42.1 43.8 35.8 25.5 16.5
$\bar{\text{IoU}}$ 27.0 24.2 20.3 10.2 3.7
Vision $\bar{P}$ 70.0 53.6 48.6 48.2 41.3
$\bar{R}$ 66.9 61.6 38.9 37.9 30.5
$\bar{\text{IoU}}$ 49.7 43.9 25.4 21.8 22.0
Vision+Audio $\bar{P}$ 56.2 51.2 37.2 33.9 21.2
$\bar{R}$ 50.6 49.0 30.9 23.4 22.0
$\bar{\text{IoU}}$ 36.9 33.4 18.4 12.1 11.1
Overall $\bar{P}$ 55.3 49.0 40.3 39.3 25.9
$\bar{R}$ 54.8 52.5 34.6 28.9 24.0
$\bar{\text{IoU}}$ 39.6 35.4 21.2 15.2 13.6
Table 3: Performance of different models on the VUE-TR benchmark across various evaluation attributes. $\bar{P}$ and $\bar{R}$ represent the AUC (Area Under Curve) values for precision and recall, respectively; while $\bar{\text{IoU}}$ denotes the AUC of intersection-over-union between prediction and ground-truth timestamp ranges, as defined in Section 6.2. GPT models are accessed via the Azure API, and Gemini is accessed via internal Google API. $\dagger$ Gemini models are evaluated by directly uploading videos. To comply with the $100$ MB upload limit, long videos are resized to a resolution of $256$ pixels. Compared to Gemini-2.0-Flash, Gemini-2.5-Pro (0325) incurs higher latency and token cost for reasoning and exhibits a higher content filtering rate, often resulting in empty outputs. $\ddagger$ GPT-4o is constrained by the Azure API’s $120$-frame limit. For videos longer than $120$ seconds, we uniformly sample $120$ frames.
@@ 7.2 Temporal Retrieval Results
Since GPT-4o API does not support audio, we extract frames at $1$ fps and feed them as input images. To adapt GPT-4o for temporal retrieval, we design a simple custom prompt with in-context examples, ensuring that the output only contains frame index ranges. An example instruction is as follows:
@@
Unlike GPT-4o, Gemini models [1] naturally support text, vision, and audio with extremely long context length, making them ideal for long video understanding. Gemini can also take raw video files as input. We evaluate two versions:
  • Gemini-2.5-Pro (0325): Offers strong reasoning and often outputs explicit Chain-of-Thought (CoT) explanations. However, when not constrained by output formatting instructions, the responses can be highly inconsistent and difficult to parse. In contrast, adding strict format requirements improves consistency but leads to degraded performance. Therefore, we choose the simple prompt without format requirement and parse the major possible formats for evaluation.
  • Gemini-2.0-Flash (stable): Provides a better output structure and lower rejection rates, making it more reliable for batch evaluations. Although slightly less capable in reasoning than Gemini-2.5-Pro, it balances performance and robustness well.
The prompt provided to Gemini models is:
Answer with time ranges and do not output explanation. What are all the time ranges corresponding to the text query: "QUERY"?
However, Gemini-0325 frequently deviates from these instructions, returning time ranges in inconsistent formats. We therefore implement careful parsing procedures to reliably extract time intervals for evaluation.
In Table 3, we report performance in multiple evaluation dimensions: video duration, query format, and query modality. Remarkably, Vidi outperforms all baseline models in all categories for the primary metric $\bar{\text{IoU}}$. In ultra-short videos, Gemini-2.0-Flash achieves the best precision, but its recall is significantly lower, leading to a lower $\bar{\text{IoU}}$. Performance across different query formats is relatively consistent, though models generally struggle more with longer and more descriptive queries. As expected, GPT-4o performs poorly in audio-only queries due to lack of audio input support. Other models perform comparably across modalities, but accuracy in audio-based queries is consistently lower than that on vision-based queries, highlighting the challenge of audio-conditioned understanding.
As shown in Figure 10, we can still observe that Vidi consistently outperforms all competitors by a significant margin across the entire range of thresholds. This performance gap even widens at higher thresholds (e.g., $\geq 0.5$), where other models degrade much more rapidly. It shows Vidi’s strength in fine-grained temporal precision, a critical capability for real-world video editing and retrieval tasks. By using more powerful backbones and dynamic token compression, Vidi-1.5 gains considerable improvement over all the metrics.
@@ 8 Related Work
In contrast, temporal retrieval requires fine-grained understanding across the full temporal span of a video instead of a few scattered glimpses. As such, temporal retrieval serves as a more rigorous benchmark for evaluating a model’s true capability in video understanding. Charades-STA [6] and DiDeMo [9] focus mainly on action grounding within short clips ($\sim 30s$). QVHighlights [14] and ActivityNet Captions [12] extend to several minutes videos, covering event descriptions and subjectively defined highlights. Recently, LongVALE [7] introduces multimodal audio-visual queries, but remains constrained to relatively short durations.
@@
Video understanding has made significant progress with the emergence of LMMs that integrate visual and textual information. Models such as the InternLM-XComposer [36, 4, 37], Qwen-VL [2, 29, 3], and LLaVA [19, 33] series have introduced innovative architectures and training strategies. InternLM-XComposer adopts a modular architecture that integrates interleaved text-image composition and comprehension with rich multilingual knowledge. Qwen-VL and its successors introduce flexible multi-image and box-conditioned input capabilities, supporting multilingual and multi-turn interaction for diverse vision-language tasks. The LLaVA series bridges visual perception with LLMs through a unified representation space, and its recent extensions, such as LLaVA-NeXT, further support video and 3D input via interleaved multimodal pretraining [17]. Although these models demonstrate impressive general-purpose reasoning, they typically operate on a limited number of frames (e.g., $\leq 64$), which restricts their effectiveness in long-form temporal retrieval.
@@
Early methods [14, 6, 9, 12] usually employ proposal-based strategies, where models first generate candidate segments and then rank them based on their semantic relevance to the query. With the emergence of video LMMs, recent research has shifted toward enhancing video-text alignment through unified frameworks. For example, TimeChat [24] connects a time-aware frame encoder and a sliding video Q-Former [18] to deal with temporal retrieval based on an instruction-tuning dataset. TimeSearch [21] proposes a temporal spotlight grounding strategy to find key events and a temporal reflection mechanism to verify time range predictions and guide the search direction. Ye et al. [34] revisit the temporal search in long-form videos by proposing a lightweight framework that reformulates temporal retrieval as a spatial search problem. Although efficient, the above methods rely on sampling only $8\sim 96$ frames, which may be insufficient to capture fine-grained temporal cues.
Despite these advancements, most existing methods still focus on short videos or rely on sparse frame sampling, and primarily limited to vision-only inputs. These constraints hinder the evaluation of models’ full multimodal reasoning abilities, particularly in understanding and retrieving moments from long, complex videos involving both visual and auditory modalities. In this work, we propose Vidi, a new approach that jointly incorporates text, vision, and audio input from hour-long videos to retrieve the most relevant temporal segments based on free-form natural language queries. Our setting pushes beyond existing benchmarks by requiring fine-grained temporal understanding across extended multimodal content.
@@ 9 Conclusion
We introduce Vidi, a large multi-modal model for real-world video understanding and editing (VUE). The first release focuses on temporal retrieval (TR), which is a foundational step in video trimming and editing, by localizing relevant segments from long videos given natural language queries. To support robust evaluation under realistic conditions, we propose VUE-TR, a new benchmark for practical video editing use cases. VUE-TR introduces key improvements over prior datasets in terms of video duration, audio support, query format, annotation quality, and metric design for multispan timestamp evaluation. Vidi leverages a unified vision-audio-language architecture with modality-aware sampling and precise token alignment. In addition, the Decomposed Attention mechanism allows the model to capture fine-grained temporal cues across diverse modalities and scales seamlessly to long videos. As demonstrated in the VUE-TR benchmark, Vidi consistently outperforms leading proprietary models such as GPT-4o and Gemini. For future work, we plan to extend Vidi to support a wider spectrum of video understanding tasks, including temporal VQA, spatio-temporal grounding, and high-level video understanding. Additionally, we aim to explore interactive editing capabilities to further bridge the gap between large models and real-world VUE applications.
@@ 10 Contributors
Chia-Wen Kuo, Dawei Du, Fan Chen, Guang Chen, Sijie Zhu, Xin Gu.
Core Contributors - Infrastructure (alphabetical order)
@@
Jiamin Yuan, Lingxi Zhang, Lu Guo, Lusha Li, Qingyu Chen, Rachel Deng, Stuart Siew, Wei Lu, Wen Zhong, Xing Mei, Xueqiong Qu, Zhenfang Chen.